In a 0.15 M NH3 solution, what percent of the NH3 is present as NH4

+? (Kb = 1.8 x 10^-5)


1
Expert's answer
2020-12-01T09:03:15-0500

Kb= a^2c. a= √(Kb/c)

a= √(1.8 x 10^-5/0.15)=0,0282842712.

W%= 0,0282842712 x 100 % = 2,82842712%.